consider the following price index in 2017:85; price index in 2018: 100; price index 2019: 110; price index in 2020: 129; price index in 2021:.151, the base year is
Answer #1

Answer

2018

=====

The price index is 100 in the base year.

Price index =( current year basket cost /base year basket cost)*100

The current year ad base year is the same so the basket cost is the same

Price index in base year =( current year basket cost /base year basket cost)*100=1*100=100